Problem 1 An electron moves in a uniform electric field, E 2 x 10° N/C from point A to point B. The distance between points A and B is 0.75m. The direction of the electric field is from point B to point A. (i) Determine the force on the electron. (ii) Determine the work done by the electric field. (i) Determine the potential difference Va - VB
Answer #1

i)

force on the electron is given by

F = qE = 1.6 x 10-19 x 2 x 106 = 3.2 x 10-13 N

ii)

workdone by the electric field

W = q x dV = q E d = 1.6 x 10-19 x 2 x 106 x 0.75 = 2.4 x 10-13 J

iii)

potential difference

V = Va - Vb = Ed = 2 x 106 x 0.75 = 1.5 x 106 v